OpenStudy (anonymous):

Find a cubic function with the given zeros. 6, -5, 2

OpenStudy (anonymous):

Here are the answer choices! (: f(x) = x^3 - 3x^2 + 28x + 60 f(x) = x^3 - 3x^2 - 28x - 60 f(x) = x^3 + 3x^2 - 28 + 60 f(x) = x^3 - 3x^2 - 28x + 60

OpenStudy (anonymous):

set the values to 0... x = 6 x -6 = 0 like that....show me what you get for the other ones

OpenStudy (anonymous):

once you get all three ...multiply all of them by each other...

OpenStudy (blurbendy):

(x-6)(x+5)(x-2) Expand this trinomial.

OpenStudy (anonymous):

I got 2x^3 - 2x^2 -30x.... that isn't one of the answer choices ):

OpenStudy (anonymous):

show me the terms that you multiplied together.....you did something wrong...you should have 4 terms as an answers

OpenStudy (anonymous):

I multiplied these together... I did the FOIL method... (x-6)(x+5)(x-2)

OpenStudy (anonymous):

you have to do the FOIL twice....but those are right

OpenStudy (anonymous):